Lillian had $15000 invested in two accounts one paying 9% simple interest and one paying 10% simple interest. how much was invested in each account it the interest after i yr is 1432?
Answer by richwmiller(17219)

We know the total amount of money invested. $15000
x+y=15000,
We know that the total interest for the year for the two accounts is $1432
0.09*x+0.1*y=1432
x=15000-y
We substitute for x
0.09*(15000-y)+0.1*y=1432
We multiply out
1350-0.09y+0.1*y=1432
We combine like terms.
0.01*y=82
Isolate y
y=8200 at 10%
x=15000-y
Calculate x
x=6800 at 9%
We check
0.09*6800+0.1*8200=1432
612+820=1432
1432=1432
Since this statement is TRUE and neither amount is negative then it is ok